# Do Whitepaper Claims Predict Market Behavior? Evidence from Cryptocurrency Factor Analysis

### v3

@@ 2.1 Cryptocurrency Narratives and Sentiment
Research on cryptocurrency narratives spans social media analysis, whitepaper studies, and sentiment measurement. Chen et al. (2019) show that Twitter sentiment predicts Bitcoin returns at short horizons, while Ante (2023) find that Elon Musk’s tweets generate significant abnormal returns for mentioned cryptocurrencies. Haykir and Yağlı (2022) document speculative bubbles driven by narrative contagion across crypto assets. Liu and Tsyvinski (2021) establish that cryptocurrency returns exhibit momentum, size effects, and exposure to market-wide factors distinct from traditional assets.
@@ 2.2 Factor Models in Cryptocurrency
Traditional asset pricing employs factor models to explain cross-sectional return variation. Fama (1970) established the theoretical foundation for efficient markets and factor-based returns. Fama and French (1993) introduced the three-factor model for equities; analogous developments in cryptocurrency have emerged more recently. Livan et al. (2011) demonstrate how random matrix theory can distinguish signal from noise in financial correlation matrices—a perspective we extend to narrative-factor comparisons. Caccioli et al. (2018) provide a comprehensive review of network-based approaches to financial systemic risk, complementing factor-based perspectives with topological analysis.
@@ 2.3 Tensor Methods in Finance
Tensor decomposition provides a natural framework for analyzing multi-way financial data. Kolda and Bader (2009) provide a comprehensive review of tensor decomposition methods, establishing the theoretical foundations for CP and Tucker decomposition. Recent work has applied these methods to financial time series: Chen et al. (2022) develop tensor factor models for high-dimensional time series, introducing TIPUP/TOPUP estimators with explicit applications to economics and finance. Wang et al. (2022) apply Tucker decomposition to high-dimensional vector autoregression, demonstrating improved performance over matrix-based approaches for multivariate time series. Han et al. (2024) develop CP factor models for dynamic tensors, providing uncorrelated latent factors directly applicable to asset pricing. Fan et al. (2013) introduce the POET estimator for high-dimensional covariance estimation with factor structure.
@@ 3.1 Market Data
We collect hourly OHLCV (open, high, low, close, volume) data via the CoinGecko API for 49 cryptocurrency assets spanning January 1, 2023 to December 31, 2024, yielding 17,543 timestamps per asset. Asset selection follows liquidity and data availability criteria, including major cryptocurrencies (BTC, ETH) and a diverse set of DeFi, infrastructure, and utility tokens.
@@ 3.2 Whitepaper Corpus
Market data (CoinGecko) 49 2-year hourly OHLCV
@@ 6.7 Limitations
  • Single data provider (CoinGecko) aggregated prices may not represent venue-specific microstructure dynamics.
@@ 7.4 Final Remarks
Funding
Supported by King’s College London resources. No external funding was received.
Conflict of Interest
@@
Ethics Approval
Not applicable.
Consent to Participate
Not applicable.
Consent for Publication
Not applicable.
Data Availability
Cryptocurrency market data obtained via CoinGecko API (publicly available). Whitepaper corpus collected from official project documentation (publicly available). Processed datasets and NLP classification outputs available at https://github.com/studiofarzulla/tensor-defi.
Code Availability
@@
Author Contribution
@@ Appendix D Whitepaper Corpus Details
Documents were obtained from official project sources, academic repositories (arXiv), and GitHub. Sources include original whitepapers (BTC, ETH, SOL, AVAX), academic papers (ADA, NEAR, GRT from arXiv), protocol specifications (ZEC, LINK), DeFi protocol documentation (AAVE, COMP, MKR, UNI), storage whitepapers (FIL, STORJ, SC, AR), and technical documentation (ICP, ARB, XMR).

### v4

# Dissensus AI Working Paper Series DAI-2508

@@ 2.1 Cryptocurrency Narratives and Sentiment
Research on cryptocurrency narratives spans social media analysis, whitepaper studies, and sentiment measurement. Chen et al. (2019) demonstrate that machine learning models trained on engineered sample dimensions can predict Bitcoin price movements, while Ante (2023) find that Elon Musk’s tweets generate significant abnormal returns for mentioned cryptocurrencies. Haykir and Yağlı (2022) document speculative bubbles driven by narrative contagion across crypto assets. Liu and Tsyvinski (2021) establish that cryptocurrency returns exhibit momentum, size effects, and exposure to market-wide factors distinct from traditional assets.
@@ 2.2 Factor Models in Cryptocurrency
Traditional asset pricing employs factor models to explain cross-sectional return variation. Fama (1970) established the theoretical foundation for efficient markets and factor-based returns. Fama and French (1993) introduced the three-factor model for equities; analogous developments in cryptocurrency have emerged more recently. Livan et al. (2011) demonstrate how random matrix theory can distinguish signal from noise in financial correlation matrices—a perspective we extend to narrative-factor comparisons. Caccioli et al. (2018) provide a comprehensive review of network-based approaches to financial systemic risk, complementing factor-based perspectives with topological analysis. More recently, Aste (2025) establishes the theoretical and generative foundations for information filtering networks, offering principled methods for extracting sparse dependency structures from high-dimensional financial data that may inform future network-aware factor models.
@@ 2.3 Tensor Methods in Finance
Tensor decomposition provides a natural framework for analyzing multi-way financial data. Kolda and Bader (2009) provide a comprehensive review of tensor decomposition methods, establishing the theoretical foundations for CP and Tucker decomposition. Recent work has applied these methods to financial time series: Chen et al. (2022) develop tensor factor models for high-dimensional time series, introducing TIPUP/TOPUP estimators with explicit applications to economics and finance. Wang et al. (2022) apply Tucker decomposition to high-dimensional vector autoregression, demonstrating improved performance over matrix-based approaches for multivariate time series. Han et al. (2024) develop CP factor models for dynamic tensors, providing uncorrelated latent factors directly applicable to asset pricing. Fan et al. (2013) introduce the POET estimator for high-dimensional covariance estimation with factor structure. Briola et al. (2026) extend PCA to incorporate graph structure, regularizing principal components with network topology—an approach that could inform future tensor decompositions where asset relationships carry meaningful structural information beyond covariance.
@@ 3.1 Market Data
We collect hourly OHLCV (open, high, low, close, volume) data via the Binance exchange API (accessed through the CCXT library) for 49 cryptocurrency assets spanning January 1, 2023 to December 31, 2024, yielding 17,543 timestamps per asset. Asset selection follows liquidity and data availability criteria, including major cryptocurrencies (BTC, ETH) and a diverse set of DeFi, infrastructure, and utility tokens.
@@ 3.2 Whitepaper Corpus
Market data (Binance) 49 2-year hourly OHLCV
@@ 6.7 Limitations
  • Single data provider (Binance) exchange prices may not represent venue-specific microstructure dynamics.
@@ 7.4 Final Remarks
Acknowledgements
Portions of this manuscript were drafted and revised with assistance from Claude (Anthropic). The author retains full responsibility for all intellectual content, analytical decisions, and interpretive claims. This work was conducted as part of the Adversarial Systems Research program at Dissensus AI. Working papers and replication materials are available through the Adversarial Systems & Complexity Research Initiative (ASCRI). Comments and correspondence are welcome at murad@dissensus.ai.
@@ Acknowledgements
Conflict of Interest.
@@
Funding.
This research received no external funding. Computational resources were provided by King’s College London.
Data Availability.
Cryptocurrency market data obtained via the Binance exchange API through CCXT (publicly available). Whitepaper corpus collected from official project documentation (publicly available). Processed datasets and NLP classification outputs available at https://github.com/studiofarzulla/tensor-defi.
Code Availability.
@@
AI Assistance.
Claude (Anthropic) assisted with manuscript drafting, code review, and statistical exposition. All research design, data analysis, and interpretation are solely the author’s.
Author Contributions.
@@ Appendix D Whitepaper Corpus Details
Documents were obtained from official project sources, academic repositories (arXiv), and GitHub. Sources include original whitepapers (BTC, ETH, SOL, AVAX), academic papers (ADA, NEAR, GRT from arXiv), protocol specifications (ZEC, LINK), DeFi protocol documentation (AAVE, COMP, MKR, UNI), storage whitepapers (FIL, STORJ, SC, AR), and technical documentation (ICP, ARB, XMR). One data quality note: the ATOM document was retrieved via a GitHub fallback mechanism and contains the Binance Smart Chain whitepaper rather than the Cosmos Hub whitepaper, meaning ATOM’s NLP classification profile reflects BSC claims. This affects 1 of 37 entities in the alignment analysis and is unlikely to alter conclusions given the overall weak alignment result.